### v1

# Benchmarking Chinese Commonsense Reasoning of LLMs: From Chinese-Specifics to Reasoning-Memorization Correlations

Abstract
We introduce CHARM, the first benchmark for comprehensively and in-depth evaluating the commonsense reasoning ability of large language models (LLMs) in Chinese, which covers both globally known and Chinese-specific commonsense. We evaluated 7 English and 12 Chinese-oriented LLMs on CHARM, employing 5 representative prompt strategies for improving LLMs’ reasoning ability, such as Chain-of-Thought. Our findings indicate that the LLM’s language orientation and the task’s domain influence the effectiveness of the prompt strategy, which enriches previous research findings. We built closely-interconnected reasoning and memorization tasks, and found that some LLMs struggle with memorizing Chinese commonsense, affecting their reasoning ability, while others show differences in reasoning despite similar memorization performance. We also evaluated the LLMs’ memorization-independent reasoning abilities and analyzed the typical errors. Our study precisely identified the LLMs’ strengths and weaknesses, providing the clear direction for optimization. It can also serve as a reference for studies in other fields. We will release CHARM at https://github.com/opendatalab/CHARM.
Benchmarking Chinese Commonsense Reasoning of LLMs: From Chinese-Specifics to Reasoning-Memorization Correlations
Jiaxing Sun1111thanks:   Equal contribution.,  Weiquan Huang222211footnotemark: 1,  Jiang Wu333311footnotemark: 1  thanks:   Project lead.,  Chenya Gu3333,  Wei Li3333,  Songyang Zhang3333,  Hang Yan3333,  Conghui He3333thanks:   Corresponding author. $1$ Wuhan University $2$ Tongji University $3$ Shanghai AI Laboratory heconghui@pjlab.org.cn
Figure 1: Construction of CHARM. CHARM encompasses both global and Chinese-specific commonsense. CHARM consists closely-interconnected reasoning and memorization tasks.
1 Introduction
Commonsense reasoning is important for the enhancement of the large language models (LLMs) Bommasani et al. (2021); Achiam et al. (2023) towards artificial general intelligence (AGI) Davis and Marcus (2015), therefore requires thorough evaluations. Numerous benchmarks evaluate the commonsense reasoning of LLMs, but most are English-based, limiting non-English evaluations Davis (2023). This paper focuses on assessing LLMs’ commonsense reasoning in a Chinese context. Currently, some commonsense reasoning benchmarks in Chinese are simply English translations Conneau et al. (2018); Ponti et al. (2020); Lin et al. (2022), which overlooks unique Chinese cultural, linguistic, regional, and historical aspects. These factors matter when Chinese users use the LLM, hence should be included in benchmarks. To effectively tackle this, we introduce CHARM, the benchmark designed to thoroughly and in-depth assess the abilities of LLMs in Chinese commonsense reasoning. It covers two domains: globally accepted commonsense (global domain) and Chinese-specific commonsense (Chinese domain). The latter includes 7 aspects: History (H), Traditional Culture and Arts (CA), Daily Life and Customs (LC), Entertainment (E), Public Figures (F), Geography (G), and Chinese Language (L). Therefore CHARM allows a thorough evaluation of LLMs’ reasoning in a Chinese context.
Prompt strategies like Chain of Thought (CoT) Wei et al. (2022) can significantly improve LLMs’ reasoning performance Wang et al. (2022, 2023b). Particularly, as the training corpus of LLMs is primarily in English Touvron et al. (2023a), studies Shi et al. (2022); Huang et al. (2023a); Zhang et al. (2023a) have shown that for non-English reasoning tasks, some LLMs perform better when reasoning in English than the native language. We evaluate 7 English and 12 Chinese-oriented LLMs on CHARM, employing 5 representative prompt strategies. The result shows that prompt strategies’ effectiveness depends on the LLMs’ orientation and the benchmark task’s domain, which enriches prior research and guides performance assessment and strategy choice for non-English LLMs.
LLMs’ commonsense reasoning relies on memorization. Exploring the correlation between memorization and reasoning offers insights into LLMs, aiding deeper understanding and suggesting ways to enhance these abilitiesBian et al. (2023). Some benchmarks Yu et al. (2023); Wang et al. (2023a); Fei et al. (2023) aid the research of memorization-reasoning relationships by incorporate tasks for assessing knowledge memorization and application (like reasoning). However, they often use the existing and disparate datasets for different tasks, resulting in a lack of intrinsic connections between these tasks. For instance, the question $Q_{rea}$ tests the LLM’s reasoning with the knowledge piece $K$. However, in memorization tasks, there probably is not any matching questions to determine if the LLM has effectively memorized $K$. Hence, if the LLM fails on $Q_{rea}$, it’s unclear whether due to poor reasoning or forgetfulness of $K$. This results in the disjointed evaluation of memorization and reasoning, failing to uncover their intrinsic links. To address this limitation, we selected suitable reasoning tasks from CHARM’s Chinese domain, and built related memorization questions for each reasoning question (see Figure 1). This design produces the closely-interconnected reasoning and memorization tasks, therefore allows for not only the concurrent evaluation of the two abilities, but also the assessment of memorization-independent reasoning, providing the clear guidance for the LLMs’ enhancement.
The contributions of this paper are as follows:
  • We present CHARM, the first benchmark for comprehensively evaluating the LLMs’ commonsense reasoning ability in Chinese, by encompassing not only the global but also the Chinese-specific commonsense.
  • We evaluated the representative prompt strategies on CHARM. Results showed that LLMs’ orientation and the task’s domain affect prompt strategy performance, which enriches previous research findings.
  • In CHARM, we built closely-interconnected reasoning and memorization tasks in Chinese commonsense domain, allowing for in-depth understanding the correlation between these abilities and precisely identifying the LLMs’ strengths and weaknesses. The design approach could serve as the reference for other fields.
2 Related Work
Commonsense Reasoning Benchmarks There are lots of English commonsense reasoning benchmarks or datasets Davis (2023), some of which have been translated to various languages, including Chinese Conneau et al. (2018); Ponti et al. (2020); Lin et al. (2022). There are also some native Chinese benchmarks related to commonsense reasoning. LogiQA Liu et al. (2020, 2023), sourced from the Chinese civil servant exam questions, evaluates the LLMs’ reading comprehension and logical reasoning across various knowledge types, including basic commonsense knowledges. CLUE Xu et al. (2020), the comprehensive Chinese language understanding benchmark, includes the Natural Language Inference (NLI) tasks, which require the LLMs’ commonsense reasoning abilities. CMMLU Li et al. (2023) is the comprehensive Chinese benchmark that encompasses exam questions from the 67 subjects. It includes tasks that test elementary commonsense and logical reasoning abilities. CORECODE Shi et al. (2023) is the dataset designed for commonsense reasoning and conflict detection in Chinese, presented in the dialogue format.
Prompt Strategy Prompt strategies such as CoT Wei et al. (2022) can effectively boost the reasoning capabilities of LLMs Wang et al. (2022, 2023b). Notably, as the LLM training corpus is primarily in English Touvron et al. (2023a), research reveals that for reasoning tasks in non-English languages, some LLMs exhibit superior performance when reasoning in English as opposed to the native language Shi et al. (2022); Zhang et al. (2023a); Huang et al. (2023a). Kim et al. (2023) proposed a novel cross-language transfer prompt method, which uses both the source and target languages to construct examples.
Benchmarks on Correlations of Memorization and Reasoning There are benchmarks which assess both the knowledge memorization and reasoning capabilities of the LLMs within specific domains. For instance, KoLA Yu et al. (2023), with its focus on world knowledge, includes tasks related to knowledge memorization and application (reasoning). SeaEval Wang et al. (2023a), emphasizing cross-language consistency and multicultural reasoning, involves tasks for cultural understanding and complex reasoning. There are also benchmarks aimed at specialized fields, like LawBench Fei et al. (2023), which include tasks for both memorization and application.
3 CHARM
CHARM is built for comprehensive and in-depth evaluation of LLMs in Chinese commonsense reasoning and revealing the intrinsic correlation between memorization and reasoning. Therefore, CHARM covers two domains, global and Chinese, using carefully selected tasks for comprehensive coverage. In addition, we chose reasoning tasks and constructed the closely-tied memorization tasks. The construction and main features of CHARM are in Figure 1. The detailed composition of CHARM is in Table 1.
3.1 Commonsense Domain
Global commonsense domain consists of universally understood commonsense. It covers objects and aspects of modern life that an individual should be aware of. It includes foundational knowledge that someone with a basic modern education is expected to know. When it involves individuals, they are globally recognized figures.
Chinese commonsense domain encompasses Chinese-specific elements. We categoried them into 7 aspects:
History (H) includes important events and figures in Chinese history, China’s dynasties, and other basic facts and shared knowledge about the history of China.
Traditional Culture and Arts (CA) encompasses Chinese traditional cultural arts, literary works, and traditional lifestyles.
Daily Life and Customs (LC) includes modern Chinese daily routines, clothing, food, housing, transportation festivals and so on.
Entertainment (E) includes the movies, television programs, music, and other entertainments in modern Chinese daily life.
Public Figures (F) encompasses the public figures well-known in Chinese society.
Geography (G) includes China’s geographical distribution, natural landscapes, and characteristic regional cultures.
Chinese Language (L) includes the fundamentals of the Chinese language, such as Chinese characters, idioms and so on.
For the two domains, especially for the above 7 aspects, we collected corresponding entities, forming the lists as shown in the Figure 1 and 5. Most of the entities are selected from Gaokao Bench [^1] Gaokao Bench is the collection of China’s university entrance exam questions, which contributes to all the 7 aspects.Zhang et al. (2023b), Douban [^2] https://www.douban.com/ is the popular user-centric cultural review platform in China, which mainly contributes to the Entertainment aspect., Hupu [^3] https://www.hupu.com/ is the large sports community popular in China, which mainly contributes to the Public Figures aspect.. Some entities are collected with the help of searching engines. We only collect the entities that are well-known in China. These entities are then used to create the commonsense reasoning questions, which belong to the corresponding domain and aspect.
Task Type Task Domain Chinese aspects Construction Question Type # Question
Reasoning Anachronisms Judgment (AJ) Chinese H, AC, LC, F [H] 2-option MCQ 150
global - [T][H] 2-option MCQ 150
Time Understanding (TU) Chinese H, AC, LC [H] 4-option MCQ 100
global - [T] 5or6-option MCQ 100
Sequence Understanding (SqU) Chinese H, CA, LC, G, L [H] 4-option MCQ 100
global - [T][H] 4-option MCQ 100
Movie and Music Recommendation (MMR) Chinese E [H] 4-option MCQ 50
global - [T] 4-option MCQ 50
Sport Understanding (SpU) Chinese F [H] 2-option MCQ 200
global - [H] 2-option MCQ 200
Natural Language Inference (NLI) Chinese G, E, L, [S][H] 3-option MCQ 100
global - [S] 3-option MCQ 100
Reading Comprehension (RC) Chinese all 7 aspects [S] 4-option MCQ 200
global - [S] 4-option MCQ 200
Memorization Anachronisms Judgment (AJ) Chinese H,AC,LC,F [H] Free-form QA 150
Time Understanding (TU) Chinese H, AC, LC [H] Free-form QA 83
Movie and Music Recommendation (MMR) Chinese E [H] Free-form QA 399
Sport Understanding (SpU) Chinese F [H] Free-form QA 127
Table 1: Overview of CHARM. The question numbers of reasoning and memorization tasks are 1800 and 759.
3.2 Reasoning Tasks
When designing the reasoning tasks in CHARM, we bear two criteria in mind. First, the tasks should span both commonsense domains, particularly the 7 Chinese aspects. Second, the global and Chinese tasks should have identical types and settings, differing only in their commonsense domains. From the existing English commonsense reasoning datasets Davis (2023); Suzgun et al. (2022), we selected the following 7 tasks:
Anachronisms Judgment (AJ) necessitates the LLM to identify anachronisms in provided sentences. This involves the LLM understanding the era associated with well-known figures, items, and events to facilitate commonsense-based reasoning. Global domain questions are the mix of translations [^4] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/anachronisms and handcrafted, while all Chinese domain questions are handcrafted.
Time Understanding (TU) requires the LLM infers a time (including year, date, moment, etc.) based on a given context, which necessitates the fundamental understanding of time-related commonsense and the capacity for mathematical reasoning. All question in the global domain are translations [^5] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/date_understanding and all in Chinese domain are handcrafted.
Sequence Understanding (SqU) requires the LLM sort a series of entities according to time or occurrence order, requiring logical reasoning based on commonsense. The global domain questions are the mix of translations [^6] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/logical_sequence and handcrafted; while all in the Chinese domain are handcrafted.
Movie and Music Recommendation (MMR) necessitates the LLM identifies the most similar matches to a variety of movies or music tracks, requiring the understanding of these popular movies and music and ability to identify their commonalities. All global domain questions are translations [^7] https://github.com/google/BIG-bench/tree/main/bigbench/benchmark_tasks/movie_recommendation, and all in the Chinese domain are handcrafted.
Sport Understanding (SpU) involves a crafted sentence with a famous athlete and a common sport action, and the LLM must assess its credibility, which demands understanding of sports and commonsense judgement. The questions in both domains are handcrafted, refering Suzgun et al. (2022) .
Natural Language Inference (NLI) gives two sentences and asks the LLM to classify their relationship as entailment, contradiction, or neutral, necessitating commonsense-based reasoning and judgement. All global domain questions are selected from CLUE Xu et al. (2020); the questions in the Chinese domain are partly from CLUE, and partly handcrafted.
Reading Comprehension (RC) gives a passage of text, and the LLM is required to reason based on it. All question in both domains are selected from LogiQA Liu et al. (2020, 2023).
The chosen tasks adequately cover both the commonsense domains, particularly the 7 aspects of the Chinese commonsense domain. This coverage enables a comprehensive assessment of LLMs’ commonsense reasoning ability in Chinese. Moreover, the Chinese-domain questions can be created following the similar types and settings as their global counterpart, facilitating the cleaner comparison of the LLMs’ performance across the domains.
All questions in the CHARM reasoning tasks are multiple-choice questions. Detailed information is in Table 1. Question examples of the tasks are in Figure 6 in Appendix B. We use regular expressions to extract the preferred choice from the generation of the LLMs Huang et al. (2023b); Li et al. (2023) and use accuracy as the metric.
3.3 Memorization Task
Shared commonsense knowledge pieces serve as a link between reasoning and memorization questions. From the 7 reasoning tasks, we chose 4 that can be readily associated in this manner, AJ, TU, MMR, SpU, refered as the Memorization-Reasoning-Interconnected (MRI) tasks, and built the reltated memorization questions.
Construction We first extract the commonsense knowledge pieces related to the entities in the corresponding reasoning questions. Information about each entity was collected to the degree sufficient to address the associated reasoning question, and then used to formulate the memorization questions. Following the Knowledge Memorization task in KoLA Yu et al. (2023), we choose free-form QA instead of multiple-choice or true/false questions, which can effectively avoid the impact of randomness. All memorization questions are handcrafted. Question examples are in Figure 7 in Appendix C. The averaged number of related memorization questions for each reasoning question are shown in Table 2.
Task AJ TU SpU MMR
Avg. # related memorization questions 2.1 3.2 2.0 8.0
Table 2: Averaged number of related memorization questions per reasoning question for each task.
Judgement and Metric For the MMR task, we use a rule-based matching method for evaluation; for the other three tasks, we use Chat-GPT for judgement. We use accuracy as the metric.
Figure 2: Distribution of CHARM construction.
3.4 Construction and Quality Assurance
The question construction of CHARM involves the following three methods:
Handcraft [H] The questions are created mannualy by the authors, based on the entities in §3.1. For reasoning questions in Chinese commonsense domain, we usually refer to the global counter-part and construct them with the same design idea and style, as shown in Figure 6.
Translate [T] We translate the existing English questions and only keep those consistent with the global commonsense domain.
Select [S] We select commonsense reasoning questions from existing Chinese datasets.
Detailed information of the construction of all tasks are shown in Table 1. The distribution of the three methods is shown in Figure 2.
We ensure the quality through these steps: First, two annotators independently verify each question, whether handcrafted or translated. Second, annotators are given only questions, not our answers, and use external resources to respond. They can also flag problematic questions. Last, a question passes if both annotators’ answers match ours; if not, it’s reviewed and modified by multiple authors.
4 Experimental Setup
4.1 Language Models
We evaluated the currently commonly used LLMs, which can be divided into two categories: (1) 7 English LLMs, including GPT series Achiam et al. (2023), LLaMA-2 Touvron et al. (2023b), and Vicuna [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5-16k and https://huggingface.co/lmsys/vicuna-13b-v1.5-16k. (2) 12 Chinese-oriented LLMs, including ChatGLM3999https://huggingface.co/THUDM/chatglm3-6b-32k, Baichuan2 Yang et al. (2023), InternLM2 Team (2023), Yi [^10] https://github.com/01-ai/Yi, DeepSeekBi et al. (2024) and Qwen Bai et al. (2023). For open-source models, we choose the chat version instead of the base version. For closed-source models, we use the official API [^11] We used the gpt-3.5-turbo-1106 version for GPT-3.5 and the gpt-4-1106-preview version for GPT-4.. Detailed information is in Table 3.
Models Open Source? Model Size Primary Language
LLaMA-2 Yes 7B, 13B, 70B English
Vicuna Yes 7B, 13B English
GPT-3.5 No undisclosed English
GPT-4 No undisclosed English
ChatGLM3 Yes 6B Chinese
Baichuan2 Yes 7B, 13B Chinese
InternLM2 Yes 7B, 20B Chinese
Yi Yes 6B, 34B Chinese
DeepSeek Yes 7B, 67B Chinese
Qwen Yes 7B, 14B, 72B Chinese
Table 3: LLMs evaluated in our experiments
LLM Chinese Commonsense Domain Global Commonsense Domain
AJ
TU
SqU
MMR
SpU
NLI
RC
Avg.
AJ
TU
SqU
MMR
SpU
NLI
RC
Avg.
GPT-3.5-1106
85.33
39.0
65.0
42.0
80.5
61.0
50.5
60.48
90.00
94.0
87.0
46.0
88.5
66.0
49.5
74.43
GPT-4-1106
96.67
60.0
85.0
74.0
86.0
77.0
62.5
77.31 95.33 98.0 97.0 66.0
90.0
72.0
72.0 84.33
LLaMA-2-7B
51.33
36.0
11.0
14.0
49.5
52.0
8.0
31.69
62.67
17.0
14.0
16.0
49.5
22.0
13.0
27.74
LLaMA-2-13B
56.00
33.0
38.0
30.0
58.0
47.0
38.0
42.86
66.67
24.0
39.0
50.0
53.5
57.0
33.5
46.24
LLaMA-2-70B
57.33
37.0
52.0
32.0
55.0
56.0
41.5
47.26
72.67
84.0
73.0
42.0
64.0
61.0
41.5
62.60
Vicuna-7B-v1.5
52.00
29.0
34.0
32.0
51.0
49.0
35.5
40.36
45.33
64.0
37.0
26.0
58.5
52.0
32.5
45.05
Vicuna-13B-v1.5
64.67
25.0
32.0
26.0
51.5
60.0
40.0
42.74
72.67
74.0
41.0
50.0
68.0
61.0
36.0
57.52
ChatGLM3-6B
66.00
40.0
59.0
38.0
77.0
72.0
37.5
55.64
34.00
69.0
71.0
28.0
75.5
63.0
34.0
53.50
Baichuan2-7B
76.00
41.0
48.0
38.0
72.0
53.0
49.5
53.93
55.33
65.0
54.0
26.0
60.5
59.0
29.0
49.83
Baichuan2-13B
85.33
40.0
48.0
46.0
72.5
66.0
51.5
58.48
77.33
74.0
58.0
40.0
71.0
61.0
39.0
60.05
InternLM2-7B
88.00
38.0
58.0
38.0
76.0
81.0
25.0
57.71
74.67
80.0
62.0
20.0
78.0
76.0
23.5
59.17
InternLM2-20B
88.00
55.0
54.0
44.0
74.5
80.0
23.0
59.79
82.67
83.0
61.0
14.0
74.5
72.0
27.0
59.17
Yi-6B
70.67
32.0
47.0
32.0
75.0
50.0
42.0
49.81
79.33
63.0
43.0
14.0
70.5
57.0
33.5
51.48
Yi-34B
96.00
55.0
89.0
76.0
88.5
72.0
51.5
75.43
88.67
92.0
87.0
56.0
89.0
70.0
47.5
75.74
DeepSeek-7B
81.33
34.0
50.0
50.0
79.5
57.0
31.5
54.76
68.00
76.0
47.0
50.0
72.5
59.0
32.5
57.86
DeepSeek-67B
96.67
57.0
83.0
92.0
87.5
77.0
34.5
75.38
90.00
95.0
86.0
22.0
88.0
73.0
39.0
70.43
Qwen-7B
70.67
38.0
55.0
48.0
71.0
57.0
49.5
55.60
74.67
78.0
69.0
50.0
72.5
55.0
36.0
62.17
Qwen-14B
87.33
54.0
77.0
60.0
82.5
66.0
55.0
68.83
84.00
83.0
83.0
44.0
84.5
71.0
40.0
69.93
Qwen-72B
98.00
59.0
91.0
84.0
86.5
84.0 67.5 81.43
94.00
92.0
93.0
64.0
93.0
71.0
63.5
81.50
Table 4: Accuracy on CHARM reasoning tasks. We selected the empirically optimal prompt strategy: XLT for English LLMs and ZH-CoT for Chinese-oriented LLMs. Bold and underline represent the first and second place respectively. Detailed results are in Table 8 and Table 9 of Appendix E.
4.2 Prompt Strategies
We selected 5 commonly used prompt strategies, and assessed the performance of the 19 LLMs on CHARM reasoning task:
Direct: The LLM does not perform intermediate reasoning and directly predicts the answer.
ZH-CoT: The LLM conducts intermediate reasoning Wei et al. (2022) in Chinese before producing the answer.
EN-CoT: The reasoning process of CoT is in English for the Chinese questionsShi et al. (2022).
Translate-EN: We use the DeepL api [^12] https://www.deepl.com/translator to translate our benchmark into English, and then use English CoT for reasoning Zhang et al. (2023a).
XLT: The template prompt Huang et al. (2023a) is used to change the original question into an English request, solve it step by step, and finally format the answer for output .
The examples for each prompt strategy are in Figure 8 in Appendix D. For all prompt strategies, we use the 3-shot setting.
5 Results and Analysis
5.1 Integrated Reasoning Performance
We show the performance of the 19 LLMs on CHARM reasoning tasks in Table 4. We only choose one representative prompt strategy: XLT for English LLMs and ZH-CoT for Chinese-oriented LLMs, which is based on our empirical conclusion in §5.2. The LLMs’ performance on the 7 aspects of the Chinese commonsense domain are shown in Table 10 in Appendix F.
Commonsense Domain We found that the LLMs exhibit inconsistent performance in the global and Chinese commonsense domains. The rankings of the English LLMs dropped in the Chinese domain compared to the global domain. For instance, GPT4 ranks first in the global domain, but in the Chinese domain, Qwen-72B outperforms all, pushing GPT4 to the second. In the Chinese domain, the performance of LLaMA-2-70B is even worse than many Chinese-oriented LLMs in the 6B-7B size range.
However, in the global domain, LLaMA-2-70B is better than all Chinese-oriented LLMs up to 20B in size, except for Qwen-14B.
5.2 Prompt Strategy Selection
Prompt Avg. all LLMs Avg. CN-LLMs Avg. EN-LLMs
Avg. Direct 46.28 48.41 42.64
all ZH-CoT 56.66 62.40 46.81
domains EN-CoT 54.46 58.19 48.06
Translate-EN 53.88 55.51 51.07
XLT 56.81 59.09 52.90
Avg. Direct 45.43 47.76 41.44
Chinese ZH-CoT 56.35 62.23 46.26
domains EN-CoT 52.06 56.36 44.68
Translate-EN 47.25 47.82 46.27
XLT 53.80 56.63 48.96
Avg. Direct 47.13 49.05 43.85
global ZH-CoT 56.96 62.57 47.35
domains EN-CoT 56.85 60.01 51.44
Translate-EN 60.50 63.20 55.87
XLT 59.82 61.56 56.84
Table 5: Averaged accuracy on CHARM reasoning tasks. “CN-LLMs” means the 12 Chinese-oriented LLMs, “EN-LLMs” means the 7 English LLMs.
We tested the combinations of the 19 LLMs and the 5 prompt strategies in CHARM reasoning tasks. Detailed results are in Table 8 and Table 9 in Appendix E. To draw some empirical conclusions, we analyze along the following two dimensions:
•  Dim1: global or Chinese commonsense domain.
•  Dim2: English or Chinese-oriented LLMs.
We average the $19\times 5$ LLM-prompt combinations along the above two dimensions, and the obtained results are in the Table 5. From the LLM dimension, it’s clear that various LLMs prefer different prompt strategies: XLT consistently excels for English LLMs among the 5 strategies, while for Chinese-oriented LLMs, despite some complexity, ZH-CoT generally performs best. From the commonsense domain dimension, strategies that use English for reasoning (like XLT, Translate-EN, etc.) are suitable for the global domain; however, ZH-CoT generally performs better in the Chinese domain.
The conclusion here differs from previous studies Shi et al. (2022); Huang et al. (2023a), which suggested that employing English for non-English reasoning tasks was more effective than using the native language. These previous studies had limitations, focusing only on English LLMs and neglecting the many Chinese-oriented LLMs developed since 2023. Furthermore, most benchmarks in these studies were merely translations from English, lacking unique cultural and linguistic characteristics in Chinese. The empirical findings with CHARM in this paper have somewhat alleviated those limitations, leading to more current and comprehensive conclusions, and of course still have the limitations, which are detailed in section Limitations.
5.3 Integrated Reasoning vs Memorization
We evaluated the correlation between the integrated reasoning and the memorization on the MRI tasks, as mentioned in §3.2. The average performance of the LLMs on the 4 MRI tasks is in Figure 3. Detailed performance on each task is in Figure 9 in Appendix G.1.
As shown in Figure 3, the 19 LLMs can be roughly divided into the three types:
•  Type I: Low memorization and low integrated reasoning ability. We found that apart from OpenAI’s GPT series, all other English LLMs belong to this type.
•  Type II: High memorization and medium integrated reasoning ability. GPT3.5 and all Chinese-oriented LLMs below 30B belong to this type. It’s worth noting that some LLMs have high memorization performance, but relatively poor integrated reasoning ability.
•  Type III: Ultra-high memorization and high integrated reasoning ability. This category includes GPT4 and the three Chinese-oriented LLMs that exceed a size of 30B.
Figure 3: Averaged accuracy across the 4 MRI tasks in the Chinese commonsense domain.
The above findings offer clear guidance for the enhancement of LLMs’ reasoning abilities in Chinese commonsense domain. For Type I, the limitation lies in the memorization. For Type II, there should be further improvement in understanding, applying knowledge, and reasoning abilities.
In addition, we also evaluated the correlation between memorization and integrated reasoning during the LLM pre-training process, details can be found in Figure 10 in Appendix G.2.
The results clearly indicate that strong memorization is the foundation of integrated reasoning. Weak memorization leads to poor reasoning, as shown by Type I LLMs. Also, factors other than memorization can cause significant differences in reasoning abilities among LLMs with similar memorization.
5.4 Memorization-Independent Reasoning
Rank Integrated Reasoning Memorization-independent Reasoning
FRMM MIB
1 DeepSeek-67B Yi-34B (↑3) GPT-4-1106 (↑2)
2 Qwen-72B DeepSeek-67B (↓1) Yi-34B (↑2)
3 GPT-4-1106 GPT-4-1106 (-) Qwen-72B (↓1)
4 Yi-34B Qwen-72B (↓2) DeepSeek-67B (↓3)
5 Qwen-14B GPT-3.5-1106 (↑2) GPT-3.5-1106 (↑2)
6 InternLM2-20B Qwen-14B (↓1) Qwen-14B (↓1)
7 GPT-3.5-1106 InternLM2-20B (↓1) InternLM2-20B (↓1)
8 InternLM2-7B InternLM2-7B (-) InternLM2-7B (-)
9 DeepSeek-7B Baichuan2-13B (↑1) Baichuan2-13B (↑1)
10 Baichuan2-13B DeepSeek-7B (↓1) DeepSeek-7B (↓1)
11 Baichuan2-7B Yi-6B (↑3) Baichuan2-7B (-)
12 ChatGLM3-6B ChatGLM3-6B(-) ChatGLM3-6B(-)
13 Qwen-7B Baichuan2-7B (↓2) Qwen-7B (-)
14 Yi-6B Qwen-7B (↓1) Yi-6B (-)
15 LLaMA-2-70B LLaMA-2-13B (↑1) LLaMA-2-13B (↑1)
16 LLaMA-2-13B LLaMA-2-70B (↓1) LLaMA-2-70B (↓1)
17 Vicuna-13B-v1.5 Vicuna-13B-v1.5 (-) Vicuna-13B-v1.5 (-)
18 Vicuna-7B-v1.5 LLaMA-2-7B (↑1) Vicuna-7B-v1.5(-)
19 LLaMA-2-7B Vicuna-7B-v1.5 (↓1) LLaMA-2-7B (-)
Table 6: Leaderboard on the MRI tasks. We propose two methods, i.e. FRMM and MIC, to compare the LLMs’ memorization-independent reasoning, as detailed in Appendix H. The arrows and numbers in brackets in the last two columns indicate changes in ranking order relative to the second column.
We propose two methods to compare the LLMs’ memorization-independent reasoning on the MRI tasks. The detail of the Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM) is in Appendix H.1, and Memorization-Independent Battles among LLMs (MIB) in Appendix H.2. The results are in Table 6. When comparing the leaderboards for integrated and memorization-independent reasoning, Type III LLMs rank at the forefront and the Type I rank at the end in all leaderboards. There is a slight variation in the ranking order within the three types of LLMs.
For the in-depth analysis, we chose Vicuna-13B, Qwen-7B, Qwen-72B as representatives for Type I, II, and III LLMs, and filtered out the reasoning questions in the MRI tasks, only keeping those with correct answers to the related memorization questions, same as the FRMM in Appendix H.1. This ensures the LLM has sufficiently memorized the commonsense knowledge required for the retained reasoning questions, thereby minimizing the impact of memorization on reasoning. There are totally 500 reasoning questions in the 4 MRI tasks, and the numbers of the retained are 106, 323 and 402 for Vicuna-13B, Qwen-7B and Qwen-72B respectively, as shown in Table 7.
Error Types If the LLMs provided incorrect answers for the retained reasoning questions, these errors can be referred to as memorization-independent reasoning errors. We conducted the manual review and analysis of their reasoning process, and classified the errors into 4 main categories.
•  Understanding Error In this case, the LLM was unable to accurately comprehend the question, including misunderstanding the content, ignoring or even modifying important information in the premise, and failing to grasp the core query of the question.
•  Knowledge Error The LLM incorporated inaccurate knowledge during the reasoning process. It’s important to highlight that the knowledge pieces related to the reasoning question were previously examined in the related memorization questions, which the LLM answered correctly. However, the LLM output incorrect information during the reasoning phase.
•  Logical Error The LLM made logical reasoning errors, such as mathematical reasoning errors, inability to reach the correct conclusion based on sufficient information, or reaching the correct conclusion but outputing the wrong option.
•  Other Errors These are other scattered, relatively rare types of errors.
We show examples of each type of errors in Figure 12 in Appendix I. The distribution of these error types are shown in Figure 4.
Discussions Obviously, the majority of errors are from logical reasoning mistakes and knowledge inaccuracies, which further provides the directions for LLMs’ enhancement. As for knowledge errors, prior studies Bian et al. (2023); Allen-Zhu and Li (2023) have indicated that the way LLMs remember and master knowledge is a relatively complex topic. Simple memorization doesn’t guarantee that LLMs can apply this knowledge accurately and skillfully during the reasoning process.
Models LLM type # Original # Retained # Incorrect
Vicuna-13B-v1.5
Type I
500 106 54
Qwen-7B
Type II
500 323 117
Qwen-72B
Type III
500 402 63
Table 7: Memorization-based filtering of reasoning questions. “Incorrect” means the incorrectly answered questions among the retained.
Figure 4: Distribution of the memorization-independent reasoning errors
6 Conclusion
This paper introduces CHARM, the first benchmark designed to comprehensively and thoroughly evaluate LLMs’ commonsense reasoning in Chinese. CHARM encompasses two counterpart commonsense domains, global and Chinese-specific, with the carefully selected tasks. We evaluated the representative prompt strategies for improving LLMs’ reasoning ability, and the empirical findings significantly enhances and supplements the conclusions of previous studies. CHARM comprises closely-interconnected reasoning and memorization tasks, helping to reveal the intrinsic correlation between memorization and reasoning of LLMs. We have evaluated the strengths and weaknesses of different LLMs and conducted the detailed analysis of memorization-independent reasoning abilities. We hope that CHARM’s approach to studying the correlations between memoriztion and reasoning can serve as a reference for similar research in other fields.
Limitations
This study conducted tests on combinations of the 19 LLMs and the 5 prompt strategies, resulting in empirical conclusions. However, many existing LLMs and prompt strategies have not yet been tested. Furthermore, the best prompt strategy for the commonsense reasoning task for the LLMs, particularly in Chinese or other non-English languages, is not static and should progress with LLM technology. This is influenced by three elements: (1) The new prompt strategies are continuously proposed, which are likely more effective. (2) The new LLMs may have different prompt strategy preference, or be less sensitive to prompt. (3) For other non-English languages with high resources, future LLMs would be continuously evolving and updating, and necessitate ongoing updates in evaluation.
The automation of the construction and evaluation of CHARM needs further improvement, including the following: (1) Most of the questions in CHARM Chinese domain are manually constructed by the author. This limits the number of benchmark questions and the range of knowledge pieces covered. (2) Regarding memorization-independent reasoning, we chose only 3 LLMs as representative and manually categorized the types of errors within CHARM. In future research, we could employ robust LLMs, like GPT4, for automated error classification and statistical analysis.
Ethical Consideration
This work involves human annotation. We have provided appropriate compensation for all annotators. The total cost of annotation for the project is about 2.2k RMB. For all annotators, we explicitly inform them about the use of the data and require them to ensure that the questions included in CHARM do not involve any social bias, ethical issues or privacy concerns during the annotation process.
Appendix A Entity and Question Examples of the 7 Chinese Commonsense Aspects
Figure 5 shows the number of questions and partial entities of each Chinese commonsense aspect we propose, as well as corresponding question examples.
Appendix B Question Examples of the Reasoning Task in CHARM
Figure 6 shows the question examples of the 7 reasoning tasks in CHARM, including both Chinese and global domains.
Appendix C Question Examples of the Memorization Task in CHARM
Figure 7 shows the questions examples of the memorization tasks in CHARM.
Appendix D Examples of Prompt Strategies
Figure 8 shows the examples of the 5 prompt strategies.
Appendix E Detailed Evaluation Results of 19 LLMs with 5 Prompt Strageties on Reasoning Tasks
We conducted a detailed evaluation of 19 different LLMs using 5 distinct prompt strategies. Table 8 and Table 9 respectively display the performance of various prompt strategies on 7 reasoning tasks in the CHARM’s Chinese commonsense domain and global commonsense domain.
Appendix F Performance of LLMs on Chinese Commonsense Knowledge Aspects
Table 10 displays the performance of LLMs in the 7 Chinese commonsense aspects. We only choose one representative prompt strategy: XLT for English LLMs and ZH-CoT for Chinese LLMs, which is based on our empirical conclusion in §5.2.
Appendix G Correlation of Memorization and Integrated Reasoning
G.1 Detailed Correlations of Memorization and Integrated Reasoning on the 4 MRI Tasks
The detailed performances of the 19 LLMs on the 4 MRI tasks are in Figure 9.
G.2 Correlation of Memorization and Integrated Reasoning throughout the LLM pretraining
We tested the intermediate checkpoint models of Baichuan2 and DeepSeek on the memorization and reasoning questions on the 4 MRI tasks. The results are shown in Figure 10.
With the increase in the number of tokens during the training process, the model’s memorization ability quickly reaches a high level (in fact, there is no particularly obvious difference between the results of the first checkpoint and the final results). This is because the knowledge involved in our task setting is the most basic commonsense, and thus widely and abundantly exists in various Chinese training corpora.
However, the improvement in reasoning performance significantly lags behind memorization. This is because to complete a reasoning task in CHARM is actually a multi-step process, requiring memorization of relevant knowledge, understanding of the question, use of knowledge for reasoning, and answering according to the requirements of the question and the demonstration of few-shot examples, etc. If an error occurs in any step of the above complex process, the reasoning task will fail.
Appendix H Leaderboard of Memorization-Independent Reasoning
It is non-trivial to acquire and compare the memorization-independent reasoning abilities of the LLMs. Intuitively, we can filter the reasoning questions by only retaining those whose related memorization questions are all correctedly answered by every LLMs. This approach ensures that each LLM has memorized the commonsense knowledge necessary for the retained reasoning questions. However, when we apply this process to all the 19 LLMs, only 28 reasoning questions remain out of the original 500 in the MRI tasks, which is obviously insufficient in number and lacks diversity, thereby introducing a high degree of uncertainty due to randomness.
Therefore, we propose two slightly more complex methods, one called Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM), the other is Memorization-Independent Battles among LLMs (MIB).
H.1 Filtering Reasoning Questions based on Mono-LLM-Memorization (FRMM)
This method is relatively simple, but has some flaws to a certain extent. For each LLM, we select reasoning questions based on its performance in memorization tasks: only retaining reasoning questions for which all related memorization questions are answered correctly. It’s clear that, after individual filtration, different LLMs will retain different reasoning questions, and even differ in the number of retained reasoning questions, as shown in the “# retained” column in Table 11.
Then, we calculate the accuracy of the retained reasoning questions for each LLM, and the results are shown in the “Retained Acc” column in Table 11. The LLMs are then ranked based on the accuracy, producing the leaderboard shown in the penultimate column of Table 6.
As mentioned above, while this method can reflect the memorization-independent reasoning abilities of LLMs to some extent, its drawback lies in that the denominator used in calculating the final ranking accuracy differs for different LLMs.
To overcome this, we have proposed the MIB method.
H.2 Memorization-Independent Battles among LLMs (MIB)
To overcome the shortcomings of the FRMM method, we refer to the pairwise battle method adopted in LLM evaluation Zheng et al. (2023). By tallying each LLM’s performance in a “round-robin” tournament of pairwise match-ups, we rank the performance of the LLMs.
Specifically, we select two LLMs at a time and filter the MRI task’s reasoning questions based on the performance of these two LLMs in memorization tasks. We only retain the reasoning questions whose related memorization questions are correctly answered by both LLMs. In this way, the two LLMs are battled under fair conditions. We then calculate the accuracy of these two LLMs on the retained reasoning questions separately, and compute the difference in accuracy as the battle score between the two models.
As shown in Figure 11, the element $E_{ij}$ represents the accuracy of LLM $i$ minus the accuracy of LLM $j$ during the battle between the two LLMs. For a total of 19 LLMs, we average each model’s scores from the 18 battles they participated in as their final score, as shown in Table 12.
Finally, we rank the LLMs based on these scores to produce the leaderboard shown in the last column of Table 6.
Appendix I Memorization-Independent Reasoning Errors
LLMs can answer memorization questions correctly, but they make mistakes when it comes to reasoning problems composed of these knowledge points. Figure 12 shows the examples of three memorization-independent reasoning errors of LLMs.
Figure 5: Entity and question examples of the commonsense aspects.
Figure 6: Examples of the reasoning tasks in CHARM.
Figure 7: Examples of the memorization tasks in CHARM.
Figure 8: Examples of prompt strategies.
LLM Prompt AJ TU SqU MMR SpU NLI RC Avg.
GPT-3.5-1106 Direct 32.00 33.0 46.0 48.0 60.0 62.0 52.5 47.64
ZH-CoT 17.33 49.0 69.0 46.0 69.0 66.0 55.0 53.05
EN-CoT 50.67 39.0 64.0 44.0 80.5 59.0 53.5 55.81
Translate-EN 87.33 44.0 54.0 40.0 76.5 64.0 49.5 59.33
XLT 85.33 39.0 65.0 42.0 80.5 61.0 50.5 60.48
GPT-4-1106 Direct 94.00 50.0 81.0 72.0 73.0 86.0 63.0 74.14
ZH-CoT 99.33 67.0 81.0 68.0 85.0 83.0 34.0 73.90
EN-CoT 98.00 65.0 79.0 70.0 83.0 83.0 50.5 75.50
Translate-EN 96.00 43.0 70.0 42.0 75.5 81.0 53.5 65.86
XLT 96.67 60.0 85.0 74.0 86.0 77.0 62.5 77.31
LLaMA-2-7B Direct 47.33 21.0 20.0 30.0 49.0 6.0 31.5 29.26
ZH-CoT 52.67 34.0 23.0 22.0 51.0 34.0 29.5 35.17
EN-CoT 44.67 34.0 20.0 4.0 50.5 40.0 26.5 31.38
Translate-EN 10.00 32.0 20.0 20.0 39.0 40.0 20.0 25.86
XLT 51.33 36.0 11.0 14.0 49.5 52.0 8.0 31.69
LLaMA-2-13B Direct 47.33 30.0 34.0 30.0 49.0 33.0 31.5 36.40
ZH-CoT 52.67 38.0 38.0 30.0 53.0 34.0 23.0 38.38
EN-CoT 52.67 34.0 38.0 2.0 49.0 1.0 35.5 30.31
Translate-EN 53.33 34.0 20.0 10.0 62.0 9.0 31.0 31.33
XLT 56.00 33.0 38.0 30.0 58.0 47.0 38.0 42.86
LLaMA-2-70B Direct 47.33 23.0 25.0 26.0 49.5 44.0 33.0 35.40
ZH-CoT 51.33 31.0 35.0 24.0 51.5 46.0 37.0 39.40
EN-CoT 55.33 27.0 31.0 26.0 59.0 56.0 41.0 42.19
Translate-EN 72.67 26.0 46.0 42.0 66.5 65.0 48.0 52.31
XLT 57.33 37.0 52.0 32.0 55.0 56.0 41.5 47.26
Vicuna-7B-v1.5 Direct 52.67 25.0 30.0 16.0 14.5 19.0 21.5 25.52
ZH-CoT 56.00 25.0 39.0 26.0 49.5 56.0 33.0 40.64
EN-CoT 53.33 28.0 26.0 18.0 40.5 55.0 40.5 37.33
Translate-EN 66.67 25.0 31.0 30.0 60.0 57.0 33.0 43.24
XLT 52.00 29.0 34.0 32.0 51.0 49.0 35.5 40.36
Vicuna-13B-v1.5 Direct 47.33 34.0 34.0 30.0 48.5 52.0 46.0 41.69
ZH-CoT 64.00 33.0 31.0 34.0 51.0 54.0 36.0 43.29
EN-CoT 62.67 30.0 32.0 24.0 50.0 50.0 33.0 40.24
Translate-EN 69.33 23.0 26.0 32.0 63.0 68.0 40.5 45.98
XLT 64.67 25.0 32.0 26.0 51.5 60.0 40.0 42.74
ChatGLM3-6B Direct 44.67 35.0 48.0 46.0 58.5 73.0 60.5 52.24
ZH-CoT 66.00 40.0 59.0 38.0 77.0 72.0 37.5 55.64
EN-CoT 55.33 36.0 58.0 36.0 76.0 75.0 32.5 52.69
Translate-EN 57.33 24.0 39.0 40.0 53.5 71.0 46.0 47.26
XLT 48.00 44.0 43.0 36.0 68.5 65.0 41.0 49.36
Baichuan2-7B Direct 44.67 31.0 37.0 24.0 59.0 35.0 56.0 40.95
ZH-CoT 76.00 41.0 48.0 38.0 72.0 53.0 49.5 53.93
EN-CoT 55.33 36.0 44.0 30.0 69.0 53.0 41.0 46.90
Translate-EN 55.33 21.0 26.0 24.0 41.5 52.0 33.5 36.19
XLT 56.00 35.0 44.0 28.0 68.0 48.0 44.0 46.14
Baichuan2-13B Direct 59.33 23.0 42.0 30.0 67.0 36.0 23.5 40.12
ZH-CoT 85.33 40.0 48.0 46.0 72.5 66.0 51.5 58.48
EN-CoT 72.00 40.0 50.0 34.0 68.0 64.0 42.5 52.93
Translate-EN 65.33 38.0 40.0 32.0 58.5 49.0 36.0 45.55
XLT 61.33 33.0 38.0 34.0 67.0 61.0 46.0 48.62
InternLM2-7B Direct 22.00 33.0 62.0 54.0 58.5 84.0 66.0 54.21
ZH-CoT 88.00 38.0 58.0 38.0 76.0 81.0 25.0 57.71
EN-CoT 77.33 42.0 59.0 38.0 73.0 78.0 38.5 57.98
Translate-EN 73.33 28.0 45.0 36.0 56.5 72.0 46.0 50.98
XLT 80.67 38.0 60.0 30.0 66.5 72.0 53.5 57.24
InternLM2-20B Direct 14.00 42.0 61.0 50.0 39.5 54.0 46.5 43.86
ZH-CoT 88.00 55.0 54.0 44.0 74.5 80.0 23.0 59.79
EN-CoT 68.67 40.0 48.0 42.0 67.0 68.0 25.0 51.24
Translate-EN 80.67 34.0 54.0 36.0 53.5 71.0 53.0 54.60
XLT 85.33 36.0 71.0 42.0 64.5 68.0 58.0 60.69
Yi-6B Direct 14.67 17.0 20.0 30.0 48.0 19.0 35.5 26.31
ZH-CoT 70.67 32.0 47.0 32.0 75.0 50.0 42.0 49.81
EN-CoT 58.67 18.0 34.0 30.0 58.0 52.0 48.5 42.74
Translate-EN 56.00 25.0 23.0 26.0 24.0 15.0 23.0 27.43
XLT 54.67 36.0 35.0 28.0 68.5 56.0 43.0 45.88
Yi-34B Direct 89.33 28.0 85.0 56.0 70.0 51.0 68.0 63.90
ZH-CoT 96.00 55.0 89.0 76.0 88.5 72.0 51.5 75.43
EN-CoT 90.00 42.0 78.0 66.0 84.5 67.0 50.0 68.21
Translate-EN 86.67 28.0 55.0 34.0 71.0 65.0 41.5 54.45
XLT 92.00 48.0 87.0 72.0 84.0 66.0 61.0 72.86
DeepSeek-7B Direct 46.67 27.0 21.0 30.0 48.0 40.0 27.5 34.31
ZH-CoT 81.33 34.0 50.0 50.0 79.5 57.0 31.5 54.76
EN-CoT 72.00 33.0 33.0 24.0 73.0 47.0 35.5 45.36
Translate-EN 68.67 18.0 28.0 36.0 59.0 72.0 40.5 46.02
XLT 55.33 32.0 39.0 36.0 51.0 37.0 35.0 40.76
DeepSeek-67B Direct 22.67 48.0 33.0 28.0 12.5 53.0 39.5 33.81
ZH-CoT 96.67 57.0 83.0 92.0 87.5 77.0 34.5 75.38
EN-CoT 84.00 53.0 73.0 58.0 82.0 73.0 35.0 65.43
Translate-EN 94.67 45.0 60.0 38.0 67.5 69.0 40.0 59.17
XLT 95.33 59.0 80.0 66.0 87.5 76.0 54.5 74.05
Qwen-7B Direct 50.67 28.0 41.0 50.0 60.5 56.0 55.5 48.81
ZH-CoT 70.67 38.0 55.0 48.0 71.0 57.0 49.5 55.60
EN-CoT 58.67 40.0 48.0 32.0 68.5 58.0 43.0 49.74
Translate-EN 62.67 23.0 31.0 26.0 60.0 54.0 45.0 43.10
XLT 62.67 26.0 47.0 40.0 63.0 50.0 50.5 48.45
Qwen-14B Direct 63.33 28.0 69.0 60.0 73.0 59.0 59.0 58.76
ZH-CoT 87.33 54.0 77.0 60.0 82.5 66.0 55.0 68.83
EN-CoT 85.33 48.0 68.0 56.0 77.5 76.0 53.0 66.26
Translate-EN 78.00 29.0 45.0 22.0 60.5 63.0 46.5 49.14
XLT 83.33 36.0 66.0 56.0 76.5 65.0 50.0 61.83
Qwen-72B Direct 90.67 36.0 85.0 78.0 80.0 84.0 77.5 75.88
ZH-CoT 98.00 59.0 91.0 84.0 86.5 84.0 67.5 81.43
EN-CoT 95.33 55.0 88.0 64.0 86.0 78.0 72.0 76.90
Translate-EN 92.00 37.0 53.0 32.0 73.0 76.0 57.0 60.00
XLT 93.33 50.0 86.0 70.0 83.0 75.0 58.5 73.69
Table 8: Accuracy of reasoning tasks in the Chinese commonsense domain of CHARM.
LLM Prompt AJ TU SqU MMR SpU NLI RC Avg.
GPT-3.5-1106 Direct 41.33 58.0 59.0 42.0 61.0 64.0 45.0 52.90
ZH-CoT 25.33 89.0 90.0 28.0 77.5 73.0 48.5 61.62
EN-CoT 59.33 85.0 84.0 42.0 86.0 68.0 50.0 67.76
Translate-EN 88.67 86.0 80.0 48.0 83.5 65.0 58.0 72.74
XLT 90.00 94.0 87.0 46.0 88.5 66.0 49.5 74.43
GPT-4-1106 Direct 90.67 83.0 92.0 70.0 88.0 78.0 74.0 82.24
ZH-CoT 92.67 100.0 90.0 34.0 88.0 76.0 50.5 75.88
EN-CoT 95.33 97.0 97.0 52.0 89.5 70.0 61.5 80.33
Translate-EN 92.67 93.0 91.0 48.0 74.0 68.0 62.0 75.52
XLT 95.33 98.0 97.0 66.0 90.0 72.0 72.0 84.33
LLaMA-2-7B Direct 43.33 20.0 20.0 28.0 51.5 18.0 20.5 28.76
ZH-CoT 51.33 18.0 22.0 26.0 50.0 31.0 22.5 31.55
EN-CoT 56.67 20.0 22.0 22.0 51.5 10.0 27.0 29.88
Translate-EN 4.67 20.0 35.0 2.0 51.5 38.0 28.0 25.60
XLT 62.67 17.0 14.0 16.0 49.5 22.0 13.0 27.74
LLaMA-2-13B Direct 53.33 21.0 35.0 24.0 51.0 35.0 20.5 34.26
ZH-CoT 54.67 15.0 35.0 14.0 51.5 32.0 27.0 32.74
EN-CoT 52.67 19.0 35.0 16.0 51.5 35.0 38.0 35.31
Translate-EN 34.67 19.0 35.0 20.0 57.0 37.0 28.0 32.95
XLT 66.67 24.0 39.0 50.0 53.5 57.0 33.5 46.24
LLaMA-2-70B Direct 48.67 33.0 33.0 30.0 50.0 58.0 20.5 39.02
ZH-CoT 46.00 20.0 35.0 32.0 51.5 38.0 35.5 36.86
EN-CoT 46.00 82.0 53.0 34.0 51.5 64.0 48.5 54.14
Translate-EN 84.67 76.0 58.0 52.0 71.0 64.0 57.5 66.17
XLT 72.67 84.0 73.0 42.0 64.0 61.0 41.5 62.60
Vicuna-7B-v1.5 Direct 15.33 22.0 33.0 12.0 49.5 22.0 16.5 24.33
ZH-CoT 52.00 50.0 50.0 10.0 50.5 53.0 31.0 42.36
EN-CoT 49.33 45.0 44.0 16.0 51.0 23.0 31.5 37.12
Translate-EN 76.00 58.0 47.0 36.0 66.5 57.0 37.5 54.00
XLT 45.33 64.0 37.0 26.0 58.5 52.0 32.5 45.05
Vicuna-13B-v1.5 Direct 55.33 58.0 39.0 28.0 48.5 59.0 30.0 45.40
ZH-CoT 71.33 71.0 49.0 14.0 53.0 62.0 33.0 50.48
EN-CoT 66.00 82.0 42.0 38.0 65.0 55.0 40.5 55.50
Translate-EN 84.00 66.0 60.0 66.0 71.0 60.0 42.0 64.14
XLT 72.67 74.0 41.0 50.0 68.0 61.0 36.0 57.52
ChatGLM3-6B Direct 44.00 33.0 57.0 42.0 63.0 80.0 38.0 51.00
ZH-CoT 34.00 69.0 71.0 28.0 75.5 63.0 34.0 53.50
EN-CoT 41.33 65.0 63.0 24.0 60.5 70.0 34.0 51.12
Translate-EN 66.67 59.0 70.0 42.0 66.5 71.0 39.5 59.24
XLT 52.67 58.0 70.0 46.0 66.0 66.0 42.5 57.31
Baichuan2-7B Direct 46.00 20.0 47.0 8.0 58.0 35.0 38.0 36.00
ZH-CoT 55.33 65.0 54.0 26.0 60.5 59.0 29.0 49.83
EN-CoT 44.00 64.0 49.0 20.0 58.5 56.0 31.5 46.14
Translate-EN 73.33 59.0 48.0 28.0 64.0 54.0 36.5 51.83
XLT 48.67 18.0 49.0 34.0 56.0 50.0 23.0 39.81
Baichuan2-13B Direct 64.00 17.0 55.0 20.0 58.0 37.0 23.5 39.21
ZH-CoT 77.33 74.0 58.0 40.0 71.0 61.0 39.0 60.05
EN-CoT 78.67 70.0 55.0 30.0 57.0 66.0 37.5 56.31
Translate-EN 73.33 68.0 51.0 36.0 61.5 61.0 42.0 56.12
XLT 70.67 75.0 49.0 42.0 69.5 61.0 31.0 56.88
InternLM2-7B Direct 46.67 61.0 65.0 46.0 67.0 79.0 53.5 59.74
ZH-CoT 74.67 80.0 62.0 20.0 78.0 76.0 23.5 59.17
EN-CoT 72.00 87.0 70.0 44.0 76.0 73.0 38.5 65.79
Translate-EN 70.67 81.0 75.0 60.0 78.0 73.0 48.5 69.45
XLT 66.00 87.0 72.0 52.0 76.5 66.0 43.5 66.14
InternLM2-20B Direct 81.33 54.0 78.0 50.0 63.5 46.0 48.0 60.12
ZH-CoT 82.67 83.0 61.0 14.0 74.5 72.0 27.0 59.17
EN-CoT 73.33 83.0 63.0 14.0 75.0 73.0 26.5 58.26
Translate-EN 82.00 84.0 89.0 40.0 76.0 68.0 46.5 69.36
XLT 87.33 84.0 80.0 70.0 79.0 70.0 47.0 73.90
Yi-6B Direct 47.33 17.0 47.0 14.0 25.0 11.0 23.0 26.33
ZH-CoT 79.33 63.0 43.0 14.0 70.5 57.0 33.5 51.48
EN-CoT 68.67 56.0 53.0 32.0 55.5 27.0 41.0 47.60
Translate-EN 72.67 57.0 62.0 32.0 69.0 37.0 35.5 52.17
XLT 54.67 44.0 60.0 62.0 70.5 59.0 41.5 55.95
Yi-34B Direct 82.67 67.0 85.0 58.0 53.5 45.0 64.0 65.02
ZH-CoT 88.67 92.0 87.0 56.0 89.0 70.0 47.5 75.74
EN-CoT 89.33 91.0 88.0 44.0 80.0 66.0 48.0 72.33
Translate-EN 78.00 85.0 83.0 48.0 76.5 64.0 54.5 69.86
XLT 88.00 88.0 86.0 70.0 93.5 60.0 58.0 77.64
DeepSeek-7B Direct 47.33 24.0 35.0 14.0 22.0 41.0 17.5 28.69
ZH-CoT 68.00 76.0 47.0 50.0 72.5 59.0 32.5 57.86
EN-CoT 75.33 74.0 40.0 16.0 53.5 47.0 35.5 48.76
Translate-EN 72.67 59.0 45.0 32.0 60.0 57.0 38.0 51.95
XLT 58.00 28.0 38.0 16.0 51.5 35.0 29.5 36.57
DeepSeek-67B Direct 37.33 83.0 18.0 2.0 39.5 49.0 37.0 37.98
ZH-CoT 90.00 95.0 86.0 22.0 88.0 73.0 39.0 70.43
EN-CoT 61.33 96.0 76.0 30.0 90.5 71.0 35.0 65.69
Translate-EN 90.67 87.0 81.0 52.0 81.0 54.0 58.5 72.02
XLT 86.00 93.0 72.0 60.0 93.0 64.0 46.0 73.43
Qwen-7B Direct 52.67 38.0 54.0 38.0 56.5 67.0 40.0 49.45
ZH-CoT 74.67 78.0 69.0 50.0 72.5 55.0 36.0 62.17
EN-CoT 74.00 81.0 65.0 36.0 73.5 66.0 35.5 61.57
Translate-EN 73.33 71.0 65.0 46.0 70.5 66.0 41.0 61.83
XLT 74.67 64.0 69.0 48.0 67.0 46.0 32.0 57.24
Qwen-14B Direct 70.00 58.0 82.0 36.0 78.0 55.0 47.5 60.93
ZH-CoT 84.00 83.0 83.0 44.0 84.5 71.0 40.0 69.93
EN-CoT 86.67 82.0 81.0 44.0 79.5 66.0 42.5 68.81
Translate-EN 86.67 72.0 85.0 48.0 78.0 64.0 48.5 68.88
XLT 80.00 79.0 83.0 48.0 79.0 65.0 45.0 68.43
Qwen-72B Direct 88.00 63.0 85.0 56.0 83.5 78.0 65.5 74.14
ZH-CoT 94.00 92.0 93.0 64.0 93.0 71.0 63.5 81.50
EN-CoT 90.00 92.0 86.0 60.0 92.5 66.0 58.0 77.79
Translate-EN 91.33 87.0 89.0 54.0 81.5 63.0 64.0 75.69
XLT 92.67 70.0 91.0 66.0 91.5 66.0 50.5 75.38
Table 9: Accuracy of reasoning tasks in the global commonsense domain of CHARM.
LLM Prompt H CA LC E F G L Avg.
GPT-3.5-1106 XLT 80.85 41.78 63.24 49.30 80.88 52.21 55.45 60.53
GPT-4-1106 XLT 92.55 58.22 86.76 73.24 86.27 71.68 67.27 76.57
LLaMA-2-7B XLT 45.74 16.44 22.06 26.76 49.02 26.55 21.82 29.77
LLaMA-2-13B XLT 51.06 38.36 39.71 36.62 57.84 46.02 30.91 42.93
LLaMA-2-70B XLT 55.85 38.36 51.47 39.44 55.39 43.36 49.09 47.57
Vicuna-7B-v1.5 XLT 49.47 32.19 33.82 38.03 50.49 44.25 32.73 40.14
Vicuna-13B-v1.5 XLT 56.91 38.36 25.00 36.62 52.45 50.44 36.36 42.31
ChatGLM3-6B ZH-CoT 64.89 37.67 54.41 49.30 75.98 53.98 48.18 54.92
Baichuan2-7B ZH-CoT 69.15 45.21 51.47 40.85 72.55 51.33 47.27 53.98
Baichuan2-13B ZH-CoT 79.26 43.84 54.41 47.89 72.55 59.29 49.09 58.05
InternLM2-7B ZH-CoT 76.60 33.56 61.76 49.30 75.49 49.56 45.45 55.96
InternLM2-20B ZH-CoT 78.72 40.41 47.06 54.93 74.02 48.67 49.09 56.13
Yi-6B ZH-CoT 61.70 42.47 54.41 38.03 75.00 44.25 36.36 50.32
Yi-34B ZH-CoT 89.36 56.16 82.35 73.24 88.73 63.72 60.91 73.50
DeepSeek-7B ZH-CoT 70.21 37.67 42.65 56.34 79.41 38.05 44.55 52.70
DeepSeek-67B ZH-CoT 87.23 55.48 75.00 87.32 86.76 52.21 52.73 70.96
Qwen-7B ZH-CoT 65.43 42.47 51.47 52.11 70.59 53.98 53.64 55.67
Qwen-14B ZH-CoT 81.91 62.33 70.59 60.56 82.84 58.41 56.36 67.57
Qwen-72B ZH-CoT 92.55 61.64 89.71 83.10 86.76 75.22 77.27 80.89
Table 10: Accuracy of reasoning questions on the 7 Chinese commonsense aspects of CHARM.
Figure 9: Accuracy of reasoning and memorization on the 4 MRI tasks.
Figure 10: Averaged accuracy of the intermediate checkpoint models throughout the LLM pretraining across the 4 MRI tasks.
LLM # Original Original Acc. # Retained Retained Acc.
DeepSeek-67B 500 84.6 409 87.04
Qwen-72B 500 84.2 402 84.33
GPT-4-1106 500 82.8 350 86.86
Yi-34B 500 82.8 355 89.86
Qwen-14B 500 76.0 331 80.97
InternLM2-20B 500 71.6 329 78.42
GPT-3.5-1106 500 69.8 226 81.86
InternLM2-7B 500 68.2 282 78.37
DeepSeek-7B 500 68.0 338 75.74
Baichuan2-13B 500 67.2 320 76.56
Baichuan2-7B 500 63.6 337 64.69
ChatGLM3-6B 500 62.4 258 65.12
Qwen-7B 500 62.0 323 63.78
Yi-6B 500 60.8 285 68.07
LLaMA-2-70B 500 49.8 107 53.27
LLaMA-2-13B 500 49.6 70 55.71
Vicuna-13B-v1.5 500 47.6 106 49.06
Vicuna-7B-v1.5 500 45.0 62 40.32
LLaMA-2-7B 500 43.8 49 46.94
Table 11: Filtering Reasoning questions based on Mono-LLM-Memorization (FRMM) on the MRI tasks.
LLM Final Score
GPT-4-1106 21.60
Yi-34B 19.52
Qwen-72B 18.43
DeepSeek-67B 18.25
GPT-3.5-1106 12.65
Qwen-14B 10.80
InternLM2-20B 10.01
InternLM2-7B 7.30
Baichuan2-13B 5.71
DeepSeek-7B 5.17
Baichuan2-7B -0.77
ChatGLM3-6B -3.11
Qwen-7B -4.04
Yi-6B -6.22
LLaMA-2-13B -15.31
LLaMA-2-70B -22.69
Vicuna-13B-v1.5 -22.92
Vicuna-7B-v1.5 -26.25
LLaMA-2-7B -28.13
Table 12: Final results of the Memorization-Independent Battles among LLMs (MIB) on the MRI tasks.
Figure 11: Results of the Memorization-Independent Battles among LLMs (MIB) on the MRI tasks.
Figure 12: Examples of the 3 types of memorization-independent reasoning errors of LLMs